3.The kinetic energy of a 20.0-N object is 5.00 × 102 J. The speed of this object is ___m/s. (Give your answer with the correct number of sig digs and do not include units).


4.A 10.0-N object is accelerated uniformly from rest at a rate of 2.5 m/s2. The kinetic energy of this object after it has accelerated a distance of 15.0 m is ___J. (Give your answer with the correct number of sig digs and do not include units).



1
Expert's answer
2021-07-19T09:50:17-0400

3)


"0.5mv^2=K\\0.5\frac{W}{g}v^2=K\\0.5\frac{20}{9.8}v^2=500\\v=22.1\frac{m}{s}"


4)


"K=mad\\K=\frac{10}{9.8}(2.5)(15)=38.3\ J\\"